14% of M&M are the brown ones that nobody wants. suppose that 7 M&M are selected at random, and let x be the number of M&M out of the 7 that are brown. use the binomial probability formula to find the probability that 2 of the M&M will be brown. Find p(x=2)
Answer #1

Solution

Given that ,

p = 0.14

q = 1 - p = 1 - 0.14=0.86

n = 7

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X = 2) = (7 C 2) * 0.142 * (0.86)7 -2

probability=0.1936